OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square.

OpenStudy (amoodarya):

x2 + 4x – 12 = 0 (x-2)^2-4-12=0 (x-2)^2-16=0 (x-2)^2=16 can you go on ?

OpenStudy (anonymous):

would x be 2, -6

OpenStudy (amoodarya):

no try it again!

OpenStudy (amoodarya):

(x-2)^2=16 (x-2)=+-4

OpenStudy (amoodarya):

x-2=4 x-2=-4

OpenStudy (anonymous):

x+2=±4 ?

OpenStudy (amoodarya):

ok x-2=4 x=6 x-2=-4 x=-2